OpenStudy (mendicant_bias):

(Introductory Real Analysis) Trying to solve the following problem, not even sure where to start:

OpenStudy (mendicant_bias):

\[\text{Given} \ x_1 > 4, \]\[x_{n+1}=\frac{1}{2}x_n+\frac{8}{x_n}\]

OpenStudy (mendicant_bias):

\[\text{a.) Use Math Induction to prove:} \ (\forall n \in N) \ x_n > 4 \]\[\text{b.) Prove directly:} \ (\forall n \in N) \ x_{n+1} < x_n\]\[\text{c.) Find} \ \lim_{n \rightarrow + \infty}x_n\] (Wait a minute, lol, I didn't even see the other prompts, I might actually understand how to figure these out, I'll leave these up here for now but otherwise take a minute to work on it.)

OpenStudy (mendicant_bias):

Alright, I'll take a crack at Part a first: The process for Math Induction, as I understand it. consists of: i.) - Prove the case for n = 1 is true. ii.) ASSUME that the case for n+1 is true. iii.) (?) Use the above to somehow justify the expression? IDR part iii.

OpenStudy (mendicant_bias):

In line 2 of my first reply, should I solve for x_n, or how should I set this up to plug in for n=1?

ganeshie8 (ganeshie8):

for part \(b\), notice that \(x_n\gt 4 \implies \dfrac{1}{x_n}\lt \dfrac{1}{4} \implies \dfrac{8}{x_n} \lt 2\) \(x_{n+1} = \dfrac{x_n}{2}+\dfrac{8}{x_n} \\~\\ \lt \dfrac{x_n}{2} + 2\\~\\ \lt \dfrac{x_n}{2} + \dfrac{x_n}{2} \\~\\ =x_n \)

OpenStudy (mendicant_bias):

I don't really think I entirely understand the substitution of 2 for 8/x_n/how it's allowed. Taking a minute to think

OpenStudy (mendicant_bias):

Yeah, I don't get it. Why are you allowed to substitute 2 for 8/x_n?

ganeshie8 (ganeshie8):

are you refering to first line ?

ganeshie8 (ganeshie8):

\[x_n\gt 4 \implies \dfrac{1}{x_n}\lt \dfrac{1}{4} \implies \dfrac{8}{x_n} \lt 2 \]

OpenStudy (mendicant_bias):

I agree that 8/x_n is less than 2, but I don't understand how it follows necessarily that \[x_{n+1} < \frac{x_n}{2}+2\]

ganeshie8 (ganeshie8):

Maybe try it like this : we know that, \(\dfrac{8}{x_n}\lt 2\) add \(\dfrac{x_n}{2}\) both sides, you get \(\dfrac{x_n}{2}+\dfrac{8}{x_n}\lt \dfrac{x_n}{2}+2\)

OpenStudy (mendicant_bias):

A statement that would make sense to me would be this: \[\frac{x_n}{2}+\frac{8}{x_n}<\frac{x_n}{2}+2\] But not relating x_n+1 to that. One minute, looking at what you said

ganeshie8 (ganeshie8):

left hand side in your recent reply is same as \(x_{n+1}\)

OpenStudy (mendicant_bias):

Alright, so knowing that x_1 is less than four from the prompt: Yeah, how the heck do I demonstrate i.) exactly? I'm used to being given an explicit formula for x_n consisting of some values an n in which I can plug in n=1 and solve two sides of an expression to make sure that it's true. How do I do it when I'm not explicitly given an expression with n?

OpenStudy (mendicant_bias):

Do all I have to do to prove it for x_1 is to just rearrange the expression we've just made to show x_1 < 4? e.g.

ganeshie8 (ganeshie8):

we cannot use part \(b\) to prove part \(a\) as we have used part \(a\) in proving part \(b\)

ganeshie8 (ganeshie8):

we must prove part \(a\) standalone

OpenStudy (mendicant_bias):

I think I'm confused about exactly what we're talking about, but nvm. In any case, the question still stands: I'm used to dealing with explicit formulas, e.g 2n <2^n, being able to plug in a value n=1 to prove that the statement is true for that case. How do I do it, conceptually, with this problem?

ganeshie8 (ganeshie8):



OpenStudy (mendicant_bias):

So i.) is just a given that x_1 > 4?

ganeshie8 (ganeshie8):

(proof by induction) Base case is already given for us : \(x_1\gt 4\) Induction hypothesis : Assume \(x_{k+1}=\dfrac{x_k}{2}+\dfrac{8}{x_k}\gt 4\) for some \(k\ge 1\). Induction step : ...

ganeshie8 (ganeshie8):

Yes, it works.

ganeshie8 (ganeshie8):

Induction step seems tricky though

OpenStudy (mendicant_bias):

Using inequalities in an induction proof is something I have almost zero grasp of, it's way more tricky to me than using an equation

ganeshie8 (ganeshie8):

we are still inducting on the index \(n\), which is the set of natural numbers. Its the statement we're trying to prove which has an inequality.

ganeshie8 (ganeshie8):

It is like proving \(2^n\gt 4\) for all \(n\gt 2\), except that in your problem the statement is a recurrence relation...

ganeshie8 (ganeshie8):

I'll try the induction step after some time... going for lunch now..

OpenStudy (mendicant_bias):

Oop, I just found some notes on this from my Prof. All I know now more is that solving this involves the Monotone Convergence Theorem.

OpenStudy (dan815):